# On the Utility of Domain-Adjacent Fine-Tuned Model Ensembles for Few-shot Problems

### v1

@@ A.7 Proof of Propositions
[…]
@@
=\;= $\frac{1}{N}\sum_{i\in\mathcal{N}}\ell(\mu_{i},y)\;\;\text{[assuming $M_{i}(X)=\mu_{i}$]}$
@@
=\;= loss of average ensemble=(DAFT-E $\text{loss of average ensemble}=\;\ell(\text{{DAFT-E}~{}\!${}^{\text{Z}}$}).$ (2)
Note that the derivation assumes $\ell$ to be convex in $\mu_{i}$, the outputs from the DAFT models, which is a common assumption for loss functions. ∎
@@
[…]
@@
$|\ell(M_{1})-\ell(M_{2})|\leq\rho(D_{\nu(1)},D_{\nu(2)}),$ (4)
[…]
@@
Table 6: Relative Improvement of DA(FT)2 compared to FT (Sentiment Analysis)
@@
Table 7: Relative Improvement of DA(FT)2 compared to FT (Text Similarity)
@@
MRPC DAFT- QQP 34.78 32.60 32.08 -6.49 24.68 18.15 20.21 10.83
DAFT- STSB 55.83 54.79 61.86 13.01 41.26 32.65 28.13 18.20
Average 45.30 43.69 46.97 3.26 32.97 25.40 24.17 14.51
QQP DAFT- MRPC 75.56 78.10 88.81 86.98 66.05 47.51 16.33 4.71
DAFT- STSB 81.24 80.61 99.21 97.98 77.47 57.41 19.36 7.38
Average 78.40 79.36 94.01 92.48 71.76 52.46 17.84 6.05
STSB DAFT- MRPC 46.89 53.02 50.07 51.96 45.97 28.01 1.26 0.64
DAFT- QQP 51.78 50.82 57.42 63.87 56.15 36.55 4.85 1.95
Average 49.33 51.92 53.74 57.91 51.06 32.28 3.05 1.30
Table 8: Relative Improvement of DA(FT)2 compared to DAFT-E (Sentiment Analysis)
@@
Table 9: Relative Improvement of DA(FT)2 compared to DAFT-E- LR (Text Similarity)
@@
MRPC DAFT- QQP -11.80 -11.25 -14.47 -13.93 -8.60 -2.85 4.15 7.20
DAFT- STSB 1.98 3.60 4.81 4.01 3.56 9.08 11.00 14.33
Average -4.91 -3.82 -4.83 -4.96 -2.52 3.12 7.57 10.77
QQP DAFT- MRPC -1.37 -1.01 1.90 -0.05 2.44 1.73 6.68 7.89
DAFT- STSB 1.82 0.39 7.51 5.83 9.49 8.55 9.46 10.65
Average 0.22 -0.31 4.70 2.89 5.96 5.14 8.07 9.27
STSB DAFT- MRPC 9.48 6.16 19.73 2.18 -2.31 -0.75 2.96 4.29
DAFT- QQP 13.12 4.63 25.59 10.19 4.50 5.87 6.61 5.64
Average 11.30 5.39 22.66 6.18 1.09 2.56 4.78 4.96
Table 10: Relative Improvement of DA(FT)2 compared to DAFT-E- RF (Text Similarity)
@@
MRPC DAFT- QQP 12.91 -2.02 -10.60 -14.67 -11.38 -7.66 -3.32 -0.07
DAFT- STSB 30.55 14.37 9.56 3.12 0.40 3.67 3.04 6.58
Average 21.73 6.17 -0.52 -5.77 -5.49 -2.00 -0.14 3.26
QQP DAFT- MRPC 5.92 2.20 2.44 -0.88 -1.49 -3.20 1.52 3.59
DAFT- STSB 9.35 3.64 8.08 4.95 5.29 3.29 4.17 6.23
Average 7.64 2.92 5.26 2.03 1.90 0.05 2.85 4.91
STSB DAFT- MRPC 11.08 5.44 -4.39 -5.82 -8.44 -6.09 -2.41 -0.63
DAFT- QQP 14.77 3.92 0.29 1.56 -2.06 0.18 1.05 0.66
Average 12.92 4.68 -2.05 -2.13 -5.25 -2.96 -0.68 0.01
(a) MRPC
(b) QQP
(c) STSB
Figure 16: Performance comparison of DAFT-E ${}^{\text{Z}}$, DAFT-E, single-best DAFT and FFT for Text Similarity task. The error interval for DAFT-E is based on the random choice of the few-shot samples used to learn the weights of the ensemble, using RF method, aggregated over 10 trials. The single best DAFT for different datasets are: MRPC (Roberta - STSB), QQP (Roberta - STSB), STSB (Roberta - MRPC).
A.8 Comparison with Model Soup
[…]
Dataset Uniform Soup DAFT-E ${}^{\text{Z}}$
Amazon 93.29% 93.72%
Cornell 85.09% 85.37%
IMDB 92.09% 92.28%
SST2 88.07% 89.55%
Tweet 86.09% 87.33%
Yelp 94.67% 95.30%
Table 12: Robustness Comparison of DAFT-E ${}^{\text{Z}}$ and Uniform Soup in terms of the set of models being combined. The results show that as we increase the number of DAFT models, the performance of both Uniform Soup and DAFT-E ${}^{\text{Z}}$ improves, with DAFT-E ${}^{\text{Z}}$ consistently outperforming Uniform Soup in the zero-shot setting. As discussed in Table 11, we are able to add many more models into DAFT-E ${}^{\text{Z}}$ than in Uniform Soup because Uniform Soup requires all models in the ensemble to have matching architectures.
DAFT models combined Uniform Soup DAFT-E ${}^{\text{Z}}$
SST2, Tweet 88.78% 89.09%
SST2, Tweet, Yelp 90.57% 90.86%
SST2, Tweet, Yelp, Cornell 91.19% 91.39%
SST2, Tweet, Yelp, Cornell, Amazon 92.09% 92.28%
Table 13: Comparison of different methods to combine pretrained models in terms of the requirements on the available (DAFT) models that are being combined and the kind of computation necessary. BPF: Method is LLM backpropagation free. SLI: Method needs to perform inference on the target task with a single LLM. NMWA No LLM model weight access is required by the method. MAE: Method can handle multi-architecture ensemble. SFPA: Method requires a single forward pass for few-shot adaptation.
This table highlights that DAFT-E ${}^{\text{Z}}$ and DAFT-E provide lightweight LLM backprop-free model combination schemes (similar to Uniform Soup and Greedy Soup), but DAFT-E ${}^{\text{Z}}$ and DAFT-E do not require access to the model weights, and can combine LLMs with different architectures. Furthermore, the few-shot adaptation in DAFT-E requires only a single forward-pass with the few-shot samples compared to Greedy Soup that requires multiple forward-passes. However, both Uniform and Greedy Soup finally need to perform inference with a single LLM while DAFT-E ${}^{\text{Z}}$ and DAFT-E have to perform inference with multiple LLMs.
Method 0-shot BPF SLI NMWA MAE SFPA
DAFT N/A N/A
Uniform Soup N/A
DAFT-E ${}^{\text{Z}}$ N/A
FT N/A
DA(FT)2 N/A
Greedy Soup
DAFT-E
A.9 Comparison with LEEP
To check the transferability of the DAFT models to a new target domain or task, we checked the LEEP scores between all the 6 datasets of the sentiment analysis (Figure 17). Interestingly, the heat map with the LEEP scores follows closely the heat map that we generated using the performance of the DAFT models (Figure 2). This bolsters the idea of DAFT models and their transferability being proportional to their performance as DAFT models. Moreover, these LEEP scores also tells us that an LEEP scored based ensemble method is possible and needs further investigation.
Figure 17: Heatmap using LEEP scores (Sentiment Analysis)

### v2

@@ A.9 Proof of Propositions
In the following, we let $\mathcal{N}=\{1,2,\cdots,N\}$ denote the set of indices of the DAFT models in the ensemble.
@@
=\;= $\frac{1}{N}\sum_{i\in\mathcal{N}}\ell(\mu_{i},y)\;\;\text{[denoting $M_{i}(X)$ as $\mu_{i}$]}$
@@
=\;= loss of average ensemble=(DAFT-E $\text{loss of average ensemble}=\;\ell(\text{{\small DAFT\!-E}~{}\!${}^{\text{\!\scriptsize Z}}$}).$ (2)
Note that the derivation assumes $\ell$ to be convex in $\mu_{i}$, the outputs from the DAFT models. We assume a convex loss function since the loss is calculated using the output probabilities of each DAFT model, which is a continuous value. Also, the most popular loss functions like MAE, or MSE are directly proportional to the distance between the predicted output $\hat{y}_{\text{pred}}$ and the true output $y$, and are convex functions (Terven et al., 2023). ∎
@@
If the ensemble weights for the $i^{th}$ model is $w_{i}$, then for the weighted ensemble method we have;
@@
Note that the base model that $M_{i}$ is built from is $B_{\kappa(i)}$. Further, $\tilde{M}_{i}$ denotes the corresponding FFT, i.e., the model built from $B_{\kappa(i)}$ by fine-tuning on the target dataset $D_{T}$. Now, let us assume that for two DAFT models $M_{1}$ and $M_{2}$ developed from the same base model and fine-tuned on different datasets, i.e., $D_{\nu(1)}$ and $D_{\nu(2)}$, to have the following bound on their losses;
$|\ell(M_{1})-\ell(M_{2})|\leq\rho(D_{\nu(1)},D_{\nu(2)}),$ (4)
where $\rho$ is an appropriately defined distance measure between the datasets $D_{\nu(1)}$ and $D_{\nu(2)}$. Then from Equation A.9 we have;
@@
Now, let us denote $\tilde{M}_{*}$ be the model fine-tuned on $D_{T}$ and performs the best among all the models when fine-tuned on $D_{T}$ (all possible FFTs). Then from 5 and by defining $\mu(M_{1},M_{2})\overset{\Delta}{=}\ell(M_{1})-\ell(M_{2})$, we have
@@
Table 9: Relative Improvement of DA(FT)2 compared to FT (Sentiment Analysis). The base model for all the models (DAFT and FT) here is Roberta-base.
@@
Table 10: Relative Improvement of DA(FT)2 compared to FT (Text Similarity). The R, B, X at the end of the DAFT names denote the base model of the DAFT models, and they denote: Roberta, Bert, Xlnet base models respectively.
@@
MRPC DAFT- QQP-R 34.78 32.60 32.08 -6.49 24.68 18.15 20.21 10.83
DAFT- STSB-R 55.83 54.79 61.86 13.01 41.26 32.65 28.13 18.20
DAFT- QQP-B 36.45 34.58 39.62 -1.60 24.19 16.03 12.41 8.69
DAFT- STSB-B 54.27 52.31 59.27 12.49 37.97 27.81 23.47 12.33
DAFT- QQP-X 33.75 29.95 35.87 -2.57 24.02 16.12 15.38 6.05
DAFT- STSB-X 54.22 53.29 59.77 13.61 40.77 30.08 25.80 16.34
Average 44.88 42.92 48.08 4.74 32.15 23.47 20.90 12.07
QQP DAFT- MRPC-R 75.56 78.10 88.81 86.98 66.05 47.51 16.33 4.71
DAFT- STSB-R 81.24 80.61 99.21 97.98 77.47 57.41 19.36 7.38
DAFT- MRPC-B 63.40 65.81 72.71 78.10 59.59 42.61 9.76 -0.62
DAFT- STSB-B 79.18 82.73 89.69 94.59 69.40 51.00 14.10 2.03
DAFT- MRPC-X 61.59 65.76 72.62 75.59 56.21 42.88 8.57 -2.41
DAFT- STSB-X 76.61 78.03 83.61 88.56 63.45 49.58 15.04 1.97
Average 72.93 75.17 84.44 86.97 65.36 48.50 13.86 2.18
STSB DAFT- MRPC-R 46.89 53.02 50.07 51.96 45.97 28.01 1.26 0.64
DAFT- QQP-R 51.78 50.82 57.42 63.87 56.15 36.55 4.85 1.95
DAFT- MRPC-B 42.88 40.53 44.46 47.67 44.77 24.74 -3.51 -3.24
DAFT- QQP-B 56.55 58.26 56.90 58.49 57.51 33.36 1.83 0.89
DAFT- MRPC-X 40.73 39.28 35.59 41.34 41.23 22.05 -5.52 -3.76
DAFT- QQP-X 51.34 53.98 52.69 55.15 53.66 30.73 0.37 -0.74
Average 48.36 49.31 49.52 53.08 49.88 29.24 -0.12 -0.71
Table 11: Relative Improvement of DA(FT)2 compared to DAFT-E (Sentiment Analysis). The base model for all the DAFT models here is Roberta-base.
@@
Table 12: Relative Improvement of DA(FT)2 compared to DAFT-E- LR (Text Similarity). The R, B, X at the end of the DAFT names denote the base model of the DAFT models, and they denote: Roberta, Bert, Xlnet base models respectively.
@@
MRPC DAFT- QQP-R -11.80 -11.25 -14.47 -13.93 -8.60 -2.85 4.15 7.20
DAFT- STSB-R 1.98 3.60 4.81 4.01 3.56 9.08 11.00 14.33
DAFT- QQP-B -10.71 -9.92 -9.59 -9.44 -8.95 -4.59 -2.61 5.13
DAFT- STSB 0.95 1.94 3.14 3.53 1.15 5.10 6.97 8.66
DAFT- QQP-X -12.47 -13.02 -12.02 -10.33 -9.08 -4.52 -0.04 2.58
DAFT- STSB-X 0.92 2.60 3.46 4.56 3.20 6.96 8.99 12.54
Average -5.19 -4.34 -4.11 -3.60 -3.12 1.53 4.74 8.41
QQP DAFT- MRPC-R -1.37 -1.01 1.90 -0.05 2.44 1.73 6.68 7.89
DAFT- STSB-R 1.82 0.39 7.51 5.83 9.49 8.55 9.46 10.65
DAFT- MRPC-B -8.20 -7.84 -6.80 -4.80 -1.55 -1.65 0.65 2.40
DAFT- STSB-B 0.66 1.57 2.37 4.01 4.51 4.13 4.64 5.13
DAFT- MRPC-X -9.22 -7.87 -6.84 -6.14 -3.63 -1.46 -0.43 0.56
DAFT- STSB-X -0.78 -1.05 -0.91 0.79 0.83 3.16 5.50 5.07
Average -2.85 -2.64 -0.46 -0.06 2.01 2.41 4.42 5.28
STSB DAFT- MRPC-R -6.23 -3.43 -3.72 -2.61 -4.53 -0.39 2.68 4.96
DAFT- QQP-R -3.11 -4.82 0.99 5.03 2.13 6.25 6.32 6.32
DAFT- MRPC-B -8.79 -11.31 -7.32 -5.35 -5.32 -2.93 -2.15 0.91
DAFT- QQP-B -0.07 -0.12 0.66 1.58 3.02 3.77 3.26 5.21
DAFT- MRPC-X -10.16 -12.10 -13.01 -9.41 -7.63 -5.03 -4.20 0.36
DAFT- QQP-X -3.39 -2.82 -2.04 -0.56 0.50 1.73 1.78 3.51
Average -5.29 -5.77 -4.07 -1.89 -1.97 0.57 1.28 3.55
Table 13: Relative Improvement of DA(FT)2 compared to DAFT-E- RF (Text Similarity). The R, B, X at the end of the DAFT names denote the base model of the DAFT models, and they denote: Roberta, Bert, Xlnet base models respectively.
@@
MRPC DAFT- QQP-R 12.91 -2.02 -10.60 -14.67 -11.38 -7.66 -3.32 -0.07
DAFT- STSB-R 30.55 14.37 9.56 3.12 0.40 3.67 3.04 6.58
DAFT- QQP-B 14.31 -0.56 -5.49 -10.21 -11.73 -9.32 -9.59 -1.99
DAFT- STSB 29.24 12.54 7.81 2.65 -1.94 -0.11 -0.70 1.29
DAFT- QQP-X 12.05 -3.98 -8.03 -11.09 -11.86 -9.25 -7.21 -4.38
DAFT- STSB-X 29.20 13.26 8.15 3.67 0.05 1.66 1.17 4.91
Average 28.02 10.39 4.53 -0.85 -3.05 -4.96 -7.17 -6.78
QQP DAFT- MRPC-R 5.92 2.20 2.44 -0.88 -1.49 -3.20 1.52 3.59
DAFT- STSB-R 9.35 3.64 8.08 4.95 5.29 3.29 4.17 6.23
DAFT- MRPC-B -1.42 -4.85 -6.29 -5.59 -5.32 -6.42 -4.21 -1.69
DAFT- STSB-B 8.10 4.86 2.92 3.15 0.50 -0.91 -0.42 0.93
DAFT- MRPC-X -2.51 -4.88 -6.34 -6.92 -7.32 -6.24 -5.25 -3.46
DAFT- STSB-X 6.56 2.16 -0.38 -0.04 -3.03 -1.84 0.40 0.88
Average 7.39 3.24 0.54 -0.83 -3.83 -4.85 -4.83 -3.99
STSB DAFT- MRPC-R 11.08 5.44 -4.39 -5.82 -8.44 -6.09 -2.41 -0.63
DAFT- QQP-R 14.77 3.92 0.29 1.56 -2.06 0.18 1.05 0.66
DAFT- MRPC-B 8.05 -3.17 -7.96 -8.48 -9.20 -8.49 -7.00 -4.47
DAFT- QQP-B 18.38 9.05 -0.04 -1.78 -1.21 -2.17 -1.86 -0.39
DAFT- MRPC-X 6.42 -4.03 -13.61 -12.40 -11.42 -10.46 -8.94 -4.99
DAFT- QQP-X 14.44 6.10 -2.72 -3.85 -3.62 -4.09 -3.26 -2.00
Average 18.46 9.18 -0.69 -3.30 -4.10 -5.72 -4.95 -5.33